OpenStudy (anonymous):

The sum of two numbers is 70. One number is 8 more than the other. What’s the smaller number?

OpenStudy (anonymous):

n1 + n2 = 70 n1 + 8 = n2 Substitute n1 + 8 + n1 = 70 2n1 = 62 n1 = 31 n2 = 39